OpenStudy (anonymous):

find the value of 3 over sqrt{3}

OpenStudy (anonymous):

\[3 \over \sqrt{3}\]

OpenStudy (ghazi):

just use \[3= \sqrt{3}*\sqrt{3}\]

OpenStudy (anonymous):

\[\sqrt{9} \]

OpenStudy (ghazi):

nope it's sqrt 3

OpenStudy (anonymous):